OpenStudy (anonymous):

Find the sum of the infinite geometric series 8+4+2+1...

hartnn (hartnn):

could you find the common ratio 'r' ?

OpenStudy (anonymous):

-2?

hartnn (hartnn):

the sum for infinite geometric series is \(\Large S_n = \dfrac{a_1}{1-r} \)

hartnn (hartnn):

how ?? you know what common ratio is, right ?

OpenStudy (anonymous):

the difference between the values?

hartnn (hartnn):

nopes, common ratio is the ratio between consecutive terms r = 2nd term/ 1st term = 3rd term / 2nd term = .... and so on

hartnn (hartnn):

like if i had 27,9,3,1,.... then r = 9/27 = 3/9 = 1/3

OpenStudy (anonymous):

so 0.33?

hartnn (hartnn):

that was just an example :P your sequence is 8+4+2+1... r = ... ?

hartnn (hartnn):

r = 4/8 = 2/4 = .... ?

OpenStudy (anonymous):

0.5? or 2?

hartnn (hartnn):

its r = 1/2 (which also = 0.5) now plug in values in the formula

hartnn (hartnn):

\(a_1\) = 1s term = 8 here

OpenStudy (anonymous):

what is a1?

OpenStudy (anonymous):

first term is also 8

hartnn (hartnn):

a1 is the notation for 1st term and in this sequence it = 8

hartnn (hartnn):

\(\Large S_n = \dfrac{a_1}{1-r} = \dfrac{8}{1-1/2} = ... ?\)

OpenStudy (anonymous):

=4?

hartnn (hartnn):

sure ?

hartnn (hartnn):

\(\dfrac{8}{1/2} = 8\times 2\)

OpenStudy (anonymous):

16!!! Sorry you are totally right! THANKYOUUU!!

hartnn (hartnn):

\(\huge \color{red}{16 ~~\checkmark \quad \ddot \smile}\)